# Differential fuzz testing to detect tampering in sensor systems and its application to arms control authentication

### v1

@@ Abstract
In future nuclear arms control treaties, it will be necessary to authenticate the hardware and software components of verification measurement systems, i.e., to ensure these systems are functioning as intended and have not been tampered with by malicious actors. While methods such as source code hashing and static analysis can help verify the integrity of software components, they may not be capable of detecting tampering with environment variables, external libraries, or the firmware and hardware of radiation measurement systems. In this article, we introduce the concept of physical differential fuzz testing as a challenge-response-style tamper indicator that can holistically and simultaneously test all the above components in a cyber-physical system. In essence, we randomly sample (or “fuzz”) the untampered system’s parameter space, including both normal and off-normal parameter values, and consider the time series of outputs as the baseline signature of the system. Re-running the same input sequence on a untampered system will produce an output sequence consistent with this baseline, while running the same input sequence on a tampered system will produce a modified output sequence and raise an alarm. We then apply this concept to authenticating the radiation measurement equipment in nuclear weapon verification systems and conduct demonstration fuzz testing measurements with a sodium iodide (NaI) gamma ray spectrometer. Because there is Poisson noise in the measured output spectra, we also use a mechanism for comparing inherently noisy or stochastic fuzzing sequences. We show that physical differential fuzz testing can detect two types of tamper attempts, and conclude that it is a promising framework for authenticating future cyber-physical systems in nuclear arms control, safeguards, and beyond.
@@ 1.1 Nuclear arms control
Nuclear arms control treaties attempt to place limits on nations’ nuclear arsenals, typically in terms of the number of deployed nuclear weapons. For instance, the New START Treaty (the only nuclear arms control agreement currently in force) limits the United States and Russia to 1550 deployed nuclear warheads each [1]. Future nuclear arms control treaties are likely to require some mechanism for directly verifying nuclear warheads, and several groups in the past decade have proposed methods for verifying the authenticity of nuclear warheads without revealing sensitive information [2, 3, 4, 5, 6, 7, 8]. Because there is a strong incentive for the weapon owner (often called the host or monitored party) to cheat the verification measurement and secretly retain their nuclear weapons or material, these proposed methods must be able to detect hoax objects and must be resistant to tampering by a malicious host. In this latter tampering problem, the term authentication is given to the process(es) by which the inspector (or monitoring party) gains confidence that the verification system functions as expected and has not been tampered with. The inspector may be an independent third party such as the International Atomic Energy Agency (IAEA), a third-party country not party to the treaty, or, as in the case of New START, another signatory of the treaty. We note that the use of the term authentication here is distinct from its usual computer security sense of validating a user’s identity. By contrast, certification is the process by which the host gains confidence in the verification system; in particular, that the system (a) does not present a safety hazard around nuclear weapons or other sensitive infrastructure, and (b) that it does not store or transmit any sensitive information. Because certification may involve intrusive access to the verification system, perhaps even without the inspector present, it is a prime opportunity for tampering by the host. Under New START, for instance, if the verification system is inspector-provided, then the inspector must send (copies of) the verification equipment to the host country for certification $30$ days in advance of an inspection, the host may partially disassemble the to-be-used equipment in an eight-hour examination (albeit with the inspector present), and the equipment is stored in a host-controlled facility between uses [9, p. 43–44].
@@
Moreover, under New START, the authentication process is rather limited, largely due to the fact that the verification measurements themselves are also rather limited by the aforementioned concerns over information security. Specifically, only simple neutron count rate verification measurements with He-3 neutron detectors paired with commercial counters are permitted, and even these are primarily for confirming the absence of nuclear material on certain delivery systems [9, p. 36]. In turn, the New START authentication procedure is primarily limited to functional testing, i.e., running basic checks that the verification system’s neutron detectors can indeed detect elevated neutron count rates from calibration sources, with little to no anti-tamper testing of the counter’s underlying micro-controller software.
@@
If direct warhead verification measurements are to be made as part of a future treaty, more stringent authentication procedures will be required, especially since the scale and complexity of the verification system will grow beyond New START levels [10, 11]. As such, recent trends in authentication have focused on simple (and joint) hardware design (e.g., [12]) and clear software design choices [13]. In the latter case, recommendations include using C rather than C++ [14], rigorously formatting and linting source code [15], and applying formal methods [16]. In addition, several field demonstrations have been carried out, including the 2022 Data Authentication Demonstration [17] and (though it focused more on verification than authentication) the 2020 LETTERPRESS exercise [18]. Despite these efforts and much other work [19, 20, 21, 22, 23, 24], authentication remains an open problem in arms control.
@@ 1.2 Fuzz random testing
Fuzz testing is a random testing technique that was first used in the late 1980’s as a simplified way of finding bugs in software programs [25, 26, 27]. Fuzz testing explores the state space of a program by probing it with random inputs called “fuzz” and observing the results, checking for unexpected results that suggest the presence of underlying bugs. In its original form, it was
@@
The key distinguishing characteristic of fuzz testing is its oracle, the mechanism for automatically checking if the program being tested produced the correct answer when given a test input. Under previous testing methods, the automated construction of a test oracle can be quite challenging. Fuzz testing reduces the oracle to a simple criterion: did the program crash or did it exit normally? Such a simple criterion allows fuzz testing to be easily and rapidly applied to a wide variety of programs. When a program crashes under fuzz testing, it is a sign that the random input drove the program into an unexpected internal state. Fuzz testing does not replace thorough functional testing [28], which utilizes inputs specified by software test cases, but it is good at exposing many coding errors.
Differential testing [29] is a technique where two implementations of a program specification are run and the outputs are compared to identify differences in behavior. These differences will typically indicate a bug in one of the implementations. In this case, a test generator with knowledge of the systems and specification can automatically create a set of inputs and the oracle is a simple comparison of the outputs of the two implementations. Differential fuzz testing [30] is a variant of differential testing where the inputs are generated randomly, further simplifying the testing process.
@@ 1.3 Overview and summary of results
This paper proposes differential fuzz testing as a tamper indicator for cyber-physical systems, with a particular application in authenticating radiation measurement equipment for arms control treaties. The high-level goal is to randomly explore the input parameter space of a reference cyber-physical system (i.e., one known not to have been tampered with), recording the output of the system. This output becomes the baseline or oracle signature. Later, when evaluating the system for tampering, the same inputs are replayed and outputs compared to the baseline. Sufficiently different outputs indicate that system has been tampered with. As such, physical differential fuzz testing can be viewed as another realization of the randomized challenge/response format that has proven useful in related nuclear verification concepts, e.g., Refs. [4, 31]. A particular advantage of our technique is that it can test the entire cyber-physical system—source code, environment variables, libraries, firmware, and hardware. A simple hash on an analysis source code file will, for instance, fail to detect tampering of an environment variable that specifies which analysis library to load.
@@ 2.1 Concept of operations (CONOPS)
We note that the second fuzz sequence (step 5) could instead (or additionally) be run before the treaty verification measurements (if permitted by the treaty) in order to guard against attacks designed to delete their own code after tampering with the verification stage. Similarly, it is important that the equipment is returned to the inspector in such a way that does not give the host an opportunity to undo their tampering, e.g., by having step 4 occur immediately after step 3. Regardless of the exact details, the key is that the second fuzz sequence occurs after the host relinquishes control over the inspector’s equipment.
A similar general CONOPS can also extend our fuzz testing approach to nuclear security/safeguards rather than arms control scenarios, though this is not our focus in the present work. For instance, in safeguards for non-nuclear-weapons states, inspection equipment is often left behind in locked cabinets at a nuclear power plant or nuclear fuel production facility, and equipment re-authentication currently depends on containment and surveillance methods (e.g., tags, seals, and video cameras). Fuzz testing could increase confidence that the containment and surveillance has not been defeated, or at least act as an additional deterrent against such tampering. In nuclear weapons states, nuclear security measures such as portal monitors in power plants or weapons facilities are also routinely employed [32, 33]. Confidence in the integrity of these systems could again be increased by fuzz testing.
[…]
@@ 2.2 Tamper detection
In traditional differential fuzz testing, the system being tested is a deterministic computer program, and thus it would suffice to check whether $P_{i}=P^{\prime}_{i}$ for all outputs $i=1,\,\ldots,\,N$. If $P_{i}\neq P^{\prime}_{i}$ for any $i$, then one concludes that the system has been tampered with. However, in our measurement process $\mathcal{M}$, we make measurements of a radioactive source, which is an inherently stochastic process. Thus, even in the absence of tampering, we expect $P_{i}\neq P^{\prime}_{i}$ in general. Instead of checking whether the sequences $\mathcal{P}$ and $\mathcal{P}^{\prime}$ are equal, then, we need to check whether they are in some sense statistically consistent with each other.
Each $P_{i}$ or $P^{\prime}_{i}$ can contain potentially arbitrary output metrics, some of which will be stochastic (e.g., the output gamma ray spectra) and some of which may be deterministic. In this work, we restrict ourselves to analyzing only the output gamma ray spectra, though we note again that the outputs could also include other useful metrics such as resource usage. As such, we essentially comparing two time series of histograms, each of which was generated from the same sequence of fuzzed input parameters. The two time series should match within Poisson statistics, holding everything else constant. Malicious code should therefore manifest via an anomaly in the time series data.
@@
where we have modified the denominator to include the expected variance in both measurements [34] and use the convention $/0\equiv 0$ 0 if $p_{ij}=p^{\prime}_{ij}=0$. Since Eq. 1 compares the bin count difference to its expected error, if $\chi^{2}/\nu\approx 1$, then the histograms $p_{i}$ and $p^{\prime}_{i}$ are said to be statistically consistent. If $\chi^{2}/\nu\ll 1$ , then $p_{i}$ and $p^{\prime}_{i}$ are much more similar than expected given the measurement statistics. Conversely if $\chi^{2}/\nu\gg 1$, then $p_{i}$ and $p^{\prime}_{i}$ differ substantially and are statistically inconsistent.
While Eq. 1 compares a single histogram pair $(p_{i},p^{\prime}_{i})$, some method is required for determining whether to alarm based on the collection of all $N$ histogram comparisons. For the purposes of the demonstrations in Section 3, we will simply set fixed $\chi^{2}/\nu$ thresholds and alarm if any histogram comparison exceeds the threshold. In general, these thresholds will depend on the magnitude of change in the reduced $\chi^{2}$ the inspector expects or wishes to be sensitive to, and setting them inherently involves balancing the risk of a false alarm against a false negative. Thus in a real scenario, the values of the thresholds should be carefully considered. In addition, as the number of measurements $N$ becomes large, the likelihood of a false positive naturally increases; a future implementation should also correct for this multiple comparison effect. Finally, we note that more advanced anomaly detection methods than the reduced $\chi^{2}$ metric could be employed, perhaps leveraging machine learning methods, but that the simpler reduced $\chi^{2}$ may be easier to negotiate into a treaty.
@@ 2.3 Demonstration radiation detection system
We note that this detector configuration is similar to the NaI detector of the Trusted Radiation Identification System (TRIS), which has historically been used to measure the gamma ray spectra of nuclear weapons [35]. However, as discussed later in Section 4, a general-purpose computer such as the Intel NUC used in this proof-of-concept is unlikely to be deployed in a real treaty verification scenario.
A radiation measurement then consists of placing a radiation source near the detector, setting configuration parameters for a pre-set dwell time, starting/stopping acquisition, and writing data to disk. A list of configuration parameters is given in Table 1. In this work, we use a ${\sim}7$ µCi Cs-137 source for both the fuzz testing measurements and the proxy treaty verification measurements. Analysis and attack codes are written in Python3 and make use of the open-source numpy [36] library.
@@
dwell time $>0$ s measurement duration (real time)
high voltage 1001001001001200120012001200 V PMT bias voltage
pulse width 0.750.750.750.752.02.02.02.0 µs detector pulse width in microseconds
fine gain 0.50.50.50.51.2×1.2$\times$1.2 × detector fine gain

### v2

@@ Abstract
In future nuclear arms control treaties, it will be necessary to authenticate the hardware and software components of verification measurement systems, i.e., to ensure these systems are functioning as intended and have not been tampered with by malicious actors. While methods such as source code hashing and static analysis can help verify the integrity of software components, they may not be capable of detecting tampering with environment variables, external libraries, or the firmware and hardware of radiation measurement systems. In this article, we introduce the concept of physical differential fuzz testing as a challenge-response-style tamper indicator that can holistically and simultaneously test all the above components in a cyber-physical system. In essence, we randomly sample (or “fuzz”) the untampered system’s parameter space, including both normal and off-normal parameter values, and consider the time series of outputs as the baseline signature of the system. Re-running the same input sequence on a untampered system will produce an output sequence consistent with this baseline, while running the same input sequence on a tampered system will produce a modified output sequence and raise an alarm. We then apply this concept to authenticating the radiation measurement equipment in nuclear weapon verification systems and conduct demonstration fuzz testing measurements with a sodium iodide (NaI) gamma ray spectrometer. Because there is Poisson noise in the measured output spectra, we also use a mechanism for comparing inherently noisy or stochastic fuzzing output sequences. We show that physical differential fuzz testing can detect two types of tamper attempts, and conclude that it is a promising framework for authenticating future cyber-physical systems in nuclear arms control, safeguards, and beyond.
@@ 1.1 Nuclear arms control
Nuclear arms control treaties attempt to place limits on nations’ nuclear arsenals, typically in terms of the number of deployed nuclear weapons. For instance, the New START Treaty (the only nuclear arms control agreement currently in force) limits the United States and Russia to 1550 deployed nuclear warheads each [newstart_nti]. Future nuclear arms control treaties are likely to require some mechanism for directly verifying nuclear warheads, and several groups in the past decade have proposed methods for verifying the authenticity of nuclear warheads without revealing sensitive information [glaser2014zero, philippe2016physical, kemp2016physical, vavrek2018experimental, hecla2018nuclear, engel2019physically, marleau2017investigation]. Because there is a strong incentive for the weapon owner (often called the host or monitored party) to cheat the verification measurement and secretly retain their nuclear weapons or material, these proposed methods must be able to detect hoax objects and must be resistant to tampering by a malicious host. In this latter tampering problem, the term authentication is given to the process(es) by which the inspector (or monitoring party) gains confidence that the verification system functions as expected and has not been tampered with. The inspector may be an independent third party such as the International Atomic Energy Agency (IAEA), a third-party country not party to the treaty, or, as in the case of New START, another signatory of the treaty. We note that the use of the term authentication here is distinct from its usual computer security sense of validating a user’s identity. By contrast, certification is the process by which the host gains confidence in the verification system; in particular, that the system (a) does not present a safety hazard around nuclear weapons or other sensitive infrastructure, and (b) that it does not allow the inspector to store or transmit any sensitive information. Because certification may involve intrusive access to the verification system, perhaps even without the inspector present, it is a prime opportunity for tampering by the host. Under New START, for instance, if the verification system is inspector-provided, then the inspector must send (copies of) the verification equipment to the host country for certification $30$ days in advance of an inspection, the host may partially disassemble the to-be-used equipment in an eight-hour examination (albeit with the inspector present), and the equipment is stored in a host-controlled facility between uses [newstart_inspection_annex, p. 43–44].
@@
Moreover, under New START, the authentication process is rather limited, largely due to the fact that the verification measurements themselves are also rather limited by the aforementioned concerns over information security. Specifically, only simple neutron count rate verification measurements with He-3 neutron detectors paired with commercial counters are permitted, and even these are primarily for confirming the absence of nuclear material on certain delivery systems [newstart_inspection_annex, p. 36]. In turn, the New START authentication procedure is primarily limited to functional testing, i.e., running basic checks that the verification system’s neutron detectors can indeed detect elevated neutron count rates from calibration sources, with little to no anti-tamper testing of the counter’s underlying micro-controller software.
@@
If direct warhead verification measurements are to be made as part of a future treaty, more stringent authentication procedures will be required, especially since the scale and complexity of the verification system will grow beyond New START levels [brotz2019trusting, hamel2018next]. As such, recent trends in authentication have focused on simple (and joint) hardware design (e.g., [kutt2019vintage]) and clear software design choices [white2001increasing]. In the latter case, recommendations include using C rather than C++ [white2005computer], rigorously formatting and linting source code [weber2023equipment], and applying formal methods [evans2015software]. In addition, several field demonstrations have been carried out, including the 2022 Data Authentication Demonstration [brotz2023data] and (though it focused more on verification than authentication) the 2020 LETTERPRESS exercise [quad2020letterpress]. Despite these efforts and much other work [tolk1997verification, kouzes2001authentication, hauck2010defining, macarthur2012simultaneous, white2008tools, white2015trends], authentication remains an open problem in arms control.
@@ 1.2 Fuzz random testing
Fuzz testing is a random testing technique that was first used in the late 1980’s as a simplified way of finding bugs in software programs [miller1990empirical, klees2018evaluating, miller2020relevance]. Fuzz testing explores the state space of a program by probing it with random inputs called “fuzz” and observing the results, checking for unexpected results that suggest the presence of underlying bugs. In its original form, it was
@@
The key distinguishing characteristic of fuzz testing is its oracle, the mechanism for automatically checking if the program being tested produced the correct answer when given a test input. Under previous testing methods, the automated construction of a test oracle can be quite challenging, and can rely on assumptions about correct system behavior. Fuzz testing reduces the oracle to a simple criterion: did the program crash or did it exit normally? Such a simple criterion allows fuzz testing to be easily and rapidly applied to a wide variety of programs. When a program crashes under fuzz testing, it is a sign that the random input drove the program into an unexpected internal state. Fuzz testing does not replace thorough functional testing [kaner1999testing], which utilizes inputs specified by software test cases, but it is good at exposing many coding errors, and does not depend on assumptions about correct system behavior.
Differential testing [mckeeman1998differential] is a technique where two implementations of a program specification are run and the outputs are compared to identify differences in behavior. These differences will typically indicate a bug in one of the implementations. In this case, a test generator with knowledge of the systems and specification can automatically create a set of inputs and the oracle is a simple comparison of the outputs of the two implementations. Differential fuzz testing [yang2011finding] is a variant of differential testing where the inputs are generated randomly, further simplifying the testing process.
@@ 1.3 Overview and summary of results
This paper proposes differential fuzz testing as a tamper indicator for cyber-physical systems, with a particular application in authenticating radiation measurement equipment for arms control treaties. The high-level goal is to randomly explore the input parameter space of a reference cyber-physical system (i.e., one known not to have been tampered with), recording the output of the system. This output becomes the baseline or oracle signature. Later, when evaluating the system for tampering, the same inputs are replayed and outputs compared to the baseline. Sufficiently different outputs indicate that system has been tampered with. In this way, we apply differential fuzz testing to the same system at two different points in time, rather than two separate implementations as in traditional differential fuzz testing. Moreover, physical differential fuzz testing can be viewed as another realization of the randomized challenge/response format that has proven useful in related nuclear verification concepts, e.g., Refs. [kemp2016physical, tobisch2023remote]. A particular advantage of our technique is that it can test the entire cyber-physical system—source code, environment variables, libraries, firmware, and hardware. A simple hash on an analysis source code file will, for instance, fail to detect tampering of an environment variable that specifies which analysis library to load.
@@ 2.1 Concept of operations (CONOPS)
We note that the second fuzz sequence (step 5) could instead (or additionally) be run before the treaty verification measurements of step 3 (if permitted by the treaty) in order to guard against attacks designed to delete their own code after tampering with the verification stage. However, this option may introduce a recertification problem, whereby the host must ensure that the inspector cannot modify the system during the second fuzz testing procedure to achieve some malicious goal of their own. Similarly, it is important that the equipment is returned to the inspector in such a way that does not give the host an opportunity to undo their tampering; this could be achieved for example by having step 4 occur immediately after step 3. Regardless of the exact details, the key is that the second fuzz sequence occurs after the host relinquishes control over the inspector’s equipment.
A similar general CONOPS can also extend our fuzz testing approach to nuclear security/safeguards rather than arms control scenarios, though this is not our focus in the present work. For instance, in safeguards for non-nuclear-weapons states, inspection equipment is often left behind in locked cabinets at a nuclear power plant or nuclear fuel production facility, and equipment re-authentication currently depends on containment and surveillance methods (e.g., tags, seals, and video cameras). Fuzz testing could increase confidence that the containment and surveillance has not been defeated, or at least act as an additional deterrent against such tampering. In nuclear weapons states, nuclear security measures such as portal monitors in power plants or weapons facilities are also routinely employed [chambers1974portal, swift2020overview]. Confidence in the integrity of these systems could again be increased by fuzz testing.
More formally, the inspector supplies some sequence of (pseudo-)random “fuzz” inputs to establish a signature of the entire measurement setup (including the radiation detector, its firmware, readout software, analysis software, libraries, and environmental variables). This sequence should be kept secret from the host. Let $I_{i}$ denote the $i\textsuperscript{th}$ set of fuzz input parameters, which could include for instance the detector calibration parameters, the system time, etc. Let $P_{i}\equiv\mathcal{M}(I_{i})$ be the corresponding set of outputs after the baseline or initial measurement process $\mathcal{M}$, and $P^{\prime}_{i}\equiv\mathcal{M}^{\prime}(I_{i})$ be the set of outputs after the test measurement process $\mathcal{M}^{\prime}$. Note that, as discussed further in Section 2.2, $\mathcal{M}$ and $\mathcal{M}^{\prime}$ include the stochastic nature of radiation detection. The $P_{i}$ and $P^{\prime}_{i}$ could include the energy spectra measured by the detector, any statistics computed from the spectra, and perhaps auxiliary metrics such as disk/network/CPU usage during the measurement. Finally, let $\mathcal{P}\equiv\{P_{1},P_{2},\ldots,P_{N}\}$ (and analogously, $\mathcal{P}^{\prime}$) denote the sequence of all outputs after making $N$ measurements with fuzz input parameter sets $I_{1},I_{2},\ldots,I_{N}$. To determine whether the system has been tampered with, we require a method for testing whether the measurement processes $\mathcal{M}$ and $\mathcal{M}^{\prime}$ differ, i.e., whether the output sequences $\mathcal{P}$ and $\mathcal{P}^{\prime}$ differ.
@@ 2.2 Tamper detection
In traditional differential fuzz testing, the system being tested is a deterministic computer program, and thus it would suffice to check whether $P_{i}=P^{\prime}_{i}$ for all outputs $i=1,\,\ldots,\,N$. If $P_{i}\neq P^{\prime}_{i}$ for any $i$, then one concludes that the system has been tampered with. However, in our measurement process $\mathcal{M}$, we make measurements of a radioactive source, which is an inherently stochastic process (both in the radioactive decay and in the detection steps). Thus, even in the absence of tampering, we expect $P_{i}\neq P^{\prime}_{i}$ in general. Instead of checking whether the sequences $\mathcal{P}$ and $\mathcal{P}^{\prime}$ are equal, then, we need to check whether they are in some sense statistically consistent with each other.
Each $P_{i}$ or $P^{\prime}_{i}$ can contain potentially arbitrary output metrics, some of which will be stochastic (e.g., the output gamma ray spectra) and some of which may be deterministic. In this work, we restrict ourselves to analyzing only the output gamma ray spectra, though we note again that the outputs could also include other useful metrics such as resource usage. As such, we essentially comparing two time series of histograms, each of which was generated from the same sequence of fuzzed input parameters. The two time series should match within Poisson statistics, holding everything else constant. Malicious code or other tampering should therefore manifest via an anomaly in the time series data.
@@
where we have modified the denominator to include the expected variance in both measurements [porter2008testing] and use the convention $0/0\equiv 0$ if $p_{ij}=p^{\prime}_{ij}=0$. Since Eq. 1 compares the bin count difference to its expected error, if $\chi^{2}/\nu\approx 1$, then the histograms $p_{i}$ and $p^{\prime}_{i}$ are said to be statistically consistent. If $\chi^{2}/\nu\ll 1$ , then $p_{i}$ and $p^{\prime}_{i}$ are much more similar than expected given the measurement statistics. Conversely if $\chi^{2}/\nu\gg 1$, then $p_{i}$ and $p^{\prime}_{i}$ differ substantially and are statistically inconsistent.
While Eq. 1 compares a single histogram pair $(p_{i},p^{\prime}_{i})$, some method is required for determining whether to alarm based on the collection of all $N$ histogram comparisons. For the purposes of the demonstrations in Section 3, we will simply set fixed $\chi^{2}/\nu$ thresholds and alarm if any histogram comparison exceeds the threshold. In general, these thresholds will depend on the magnitude of change in the reduced $\chi^{2}$ the inspector expects or wishes to be sensitive to, and setting them inherently involves balancing the risk of a false alarm against a false negative. Thus in a real scenario, the values of the thresholds should be carefully considered. In addition, as the number of measurements $N$ becomes large, the likelihood of having at least one false positive naturally increases; a future implementation should also correct for this multiple comparison effect. Finally, we note that more advanced anomaly detection methods than the reduced $\chi^{2}$ metric could be employed, perhaps leveraging machine learning methods, but that the simpler reduced $\chi^{2}$ may be easier to negotiate into a treaty.
@@ 2.3 Demonstration radiation detection system
We note that this detector configuration is similar to the NaI detector of the Trusted Radiation Identification System (TRIS), which has historically been used to measure the gamma ray spectra of nuclear weapons [seager2001trusted]. However, as discussed later in Section 4, a general-purpose computer such as the Intel NUC used in this proof-of-concept is unlikely to be deployed in a real treaty verification scenario.
A radiation measurement then consists of placing a radiation source near the detector, setting configuration parameters for a pre-set dwell time, starting/stopping acquisition, and writing data to disk. A list of configuration parameters is given in Table 1. In this work, we use a ${\sim}7$ µCi Cs-137 source approximately $30$ cm from the detector face for both the fuzz testing measurements and the proxy treaty verification measurements. As discussed in Section 4, in future implementations this source-to-detector distance could also be fuzzed to provide an additional layer of security. Analysis and attack codes are written in Python3 and make use of the open-source numpy [harris2020array] library.
@@
dwell time >0>0 s measurement duration (real time)
high voltage 10010012001200 V PMT bias voltage
pulse width 0.750.752.02.0 µs detector pulse width in microseconds
[…]